OpenStudy (anonymous):

Item retailing for $4,500.00 and is marked up 100%, what is the cost and how does one arrive at the answer?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

100% is just the price that it already is., so if it is marked UP 100%, you just add the price that it already was, so $4,500+100%=$4,500+$4,500=$9,000

OpenStudy (anonymous):

good?

OpenStudy (anonymous):

No, but thank you. The selling price is $4,500. The store marked up the item by 100% - I need to find what the store paid for the item

OpenStudy (anonymous):

I don't think the store received the item for $0.00

OpenStudy (anonymous):

haha oh ok. then you just divide by 2, because then 4,500 would be 200%, so 100%=2,250

OpenStudy (anonymous):

Why by 2..I'm so confused?????

OpenStudy (anonymous):

I need to show work for a class....so basically from what I'm getting is The rate of selling price is 100% and the rate of markup is 100% so you add the two together for a rate of cost = 200% is that correct?

OpenStudy (anonymous):

because if they marked it up by 100%, then that means that 4,500=200%. therefore, 100% is 4,500/2, which is the price they originally got. it's easier to explain the other way. A store buys this product for $2,250, and want to sell it for a major profit, so they increase the price by 100%. if 100% is the price that they bought it for, then 2,250+100%=2,250+2,250=4,500.

OpenStudy (anonymous):

Okay - thank you sooooooooooooooooooooooooo much. Have a great day.

OpenStudy (anonymous):

haha can I have a medal??

OpenStudy (anonymous):

ABSOLUTELY - if you tell me how to do it...first time on here.

OpenStudy (anonymous):

next to any of my answers there should be a blue button that says "good answer". that's it!

OpenStudy (anonymous):

Bloody hell ...no blue buttom w/good answer...now what? Trying to press the picture of the medal thingy and no luck.

OpenStudy (anonymous):

to the right and directly above my answer? do you see where it says "zero medals for this answerer"?